OpenStudy (anonymous):

The legs of a right triangle measure 3x and 15. If the hypotenuse measures 3x + 3, what is the value of x?

OpenStudy (anonymous):

\[( 3x - 3 ) ^2 = ( 3x )^2 + ( 15 )^2 \\ 9x^2 - 18x + 9 = 9x^2 + 225 \\ 18x = 215 \\ x = 12 \\ \]